If you go on a road trip of 210 miles in the mountains and 8/15 of the trip is downhill, how many miles of the trip are not downhill? Reduce to the lowest term
Answer by macston(5194)

.
If 8/15 is downhill, 7/15 is not downhill.
.
(7/15)(210 miles)=1470/15 miles=98 miles